# Quantum and Classical Communication Complexity of Permutation-Invariant Functions

### v1

@@ 3.1 Quantum Communication Complexity Lower Bound
In this section, our goal is to obtain a lower bound on the quantum communication complexity for permutation-invariant functions (Theorem 3.2). Towards this end, we show that every permutation-invariant function $f$ can be reduced to $\mathsf{ESetInc}$ (defined in Definition 2.3) and exhibit a lower bound for $\mathsf{ESetInc}$ (Lemma 3.4). Additionally, Lemma 3.4 implies if $|x|=a,|y|=b$, then the cost to distinguish $|x\land y|=c-g$ from $|x\land y|=c+g$ is related to the smallest and the second smallest number in ${[a-c,c,b-c,n-a-b+c]}$.
@@
Fix $n,a,b\in\mathbb{Z}$. Consider $c$ and $g$ such that $c+g,c-g\in\mathbb{Z}$. Let $n_{1}\coloneqq\min\{[a-c,c,b-c,n-a-b+c]\}$ and $n_{2}\coloneqq\min\left(\{[a-c,c,b-c,n-a-b+c]\}\setminus\left\{n_{1}\right\}\right)$. We have
$Q^{*}\left(\mathsf{ESetInc}_{a,b,c,g}^{n}\right)=\Omega\left(\frac{\sqrt{n_{1}n_{2}}}{g}\right).$
@@
By the definitions of $f_{a,b}$ and jump of $f_{a,b}$, any quantum protocol computing $f$ can also compute $\mathsf{ESetInc}^{n}_{a,b,c,g}$ for any $a,b$ and any jump $(c,g)\in\mathscr{J}(f_{a,b})$. Therefore, given a jump $(c,g)$ for $f_{a,b}$, the cost of computing $\mathsf{ESetInc}^{n}_{a,b,c,g}$ lower bounds the cost of computing $f$. By Lemma 3.4, we have $Q^{*}(f)\geq\frac{\sqrt{n_{1}n_{2}}}{g}$ for any jump $(c,g)$ in $f_{a,b}$, where $n_{1},n_{2}$ are the smallest and the second smallest number in $\{[a-c,c,b-c,n-a-b+c]\}$. We conclude that $Q^{*}(f)=\Omega\left(m(f)\right)$ as desired. ∎
@@
Fix $n,a,b\in\mathbb{Z}$. Consider $c$ and $g$ such that $c+g,c-g\in\mathbb{Z}$. Let $n_{1}\coloneqq\min\{[a-c,c,b-c,n-a-b+c]\}$ and $n_{2}\coloneqq\min\left(\{[a-c,c,b-c,n-a-b+c]\}\setminus\left\{n_{1}\right\}\right)$. We have
$Q^{*}\left(\mathsf{ESetInc}_{a,b,c,g}^{n}\right)\geq Q^{*}\left(\mathsf{ESetInc}_{n_{1}+n_{2},n_{1}+n_{2},n_{1},g}^{n_{1}+3n_{2}}\right).$
@@
For $n_{1},n_{2}\in\mathbb{N}$ such that $n_{1}\leq n_{2}$, we have
$Q^{*}\left(\mathsf{ESetInc}_{n_{1}+n_{2},n_{1}+n_{2},n_{1},g}^{n_{1}+3n_{2}}\right)=\Omega\left(\frac{\sqrt{n_{1}n_{2}}}{g}\right).$
We use the following two results on $\mathsf{ESetInc}$ to show Lemmas 3.5 and 3.6. Specifically, Lemma 3.7 is a variant of Lemma 4.1 in [19] and shows some reduction methods to the instances of the Exact Set-Inclusion Problem. Lemma 3.8 is a generalization of Theorem 5 in [9] proved by pattern matrix method and shows the lower bound of a special instance of the Exact Set-Inclusion Problem. The proofs of Lemmas 3.7 and 3.8 are given in Appendix B.1.
@@
Fix $n,a,b\in\mathbb{Z}$. Consider $c$ and $g$ such that $c+g,c-g\in\mathbb{Z}$. The following relations hold.
@@
    $Q^{*}\left(\mathsf{ESetInc}_{a,b,c,g}^{n}\right)\leq Q^{*}\left(\mathsf{ESetInc}_{a+\ell_{1}+\ell_{3},b+\ell_{2}+\ell_{3},c+\ell_{3},g}^{n+\ell}\right)$ for integers $\ell_{1},\ell_{2},\ell_{3}\geq 0$ such that $\ell_{1}+\ell_{2}+\ell_{3}\leq\ell$;
@@
    $Q^{*}\left(\mathsf{ESetInc}_{a,b,c,g}^{n}\right)=Q^{*}\left(\mathsf{ESetInc}_{a,n-b,a-c,g}^{n}\right)=Q^{*}\left(\mathsf{ESetInc}_{n-a,b,b-c,g}^{n}\right)$;
@@
    $Q^{*}\left(\mathsf{ESetInc}_{a,b,c,g}^{n}\right)\leq Q^{*}\left(\mathsf{ESetInc}_{ka,kb,kc,kg}^{kn}\right)$, where $k\geq 1$ is an integer.
@@
For every $k\in\mathbb{Z}$, if $l$ is a half-integer and $l\leq k/2$, then $Q^{*}\left(\mathsf{ESetInc}^{4k}_{2k,k,l,1/2}\right)=\Omega\left(\sqrt{kl}\right)$.
@@
$Q^{*}\left(\mathsf{ESetInc}^{n_{1}+3n_{2}}_{n_{1}+n_{2},n_{1}+n_{2},n_{1},g}\right)$ $=Q^{*}\left(\mathsf{ESetInc}^{n_{1}+n_{2}+n_{2}+n_{2}}_{n_{1}+n_{2},n_{1}+n_{2},n_{1},g}\right)$
$\leq Q^{*}\left(\mathsf{ESetInc}^{n_{1}+n_{2}+n_{3}+n_{4}}_{n_{1}+n_{2},n_{1}+n_{3},n_{1},g}\right)$
$=Q^{*}\left(\mathsf{ESetInc}^{n}_{a,b,c,g}\right).$
If $n_{2}=b-c$ or $n-a-b+c$, the argument is similar. ∎
@@
$Q^{*}\left(\mathsf{ESetInc}^{n_{1}+3n_{2}}_{n_{1}+n_{2},n_{1}+n_{2},n_{1},g}\right)\geq Q^{*}\left(\mathsf{ESetInc}^{m_{1}+3m_{2}}_{m_{1}+m_{2},m_{1}+m_{2},m_{1},1/2}\right).$
@@
    $Q^{*}\left(\mathsf{ESetInc}^{m_{1}+3m_{2}}_{m_{1}+m_{2},m_{1}+m_{2},m_{1},1/2}\right)=\Omega\left(1\right)=\Omega\left(\sqrt{m_{1}m_{2}}\right).$
@@
    By LABEL:{le:SecInc} and 3.8, we have
    $Q^{*}\left(\mathsf{ESetInc}^{m_{1}+3m_{2}}_{m_{1}+m_{2},m_{1}+m_{2},m_{1},1/2}\right)$ $=Q^{*}\left(\mathsf{ESetInc}^{4m^{\prime}_{2}+l}_{2m^{\prime}_{2}+l_{1},m^{\prime}_{2}+l_{2},m_{1},1/2}\right)$
    $\geq Q^{*}\left(\mathsf{ESetInc}^{4m^{\prime}_{2}}_{2m^{\prime}_{2},m^{\prime}_{2},m_{1},1/2}\right)$
@@
    $k\leq\frac{1}{2}\cdot\left\lfloor\frac{2m_{1}}{3}\right\rfloor\leq\frac{1}{2}\cdot\left\lfloor\frac{m_{1}}{6}+\frac{m_{2}}{2}\right\rfloor\leq\frac{m}{2}\enspace,$ (2)
@@
    $l-\left(l_{1}+l_{2}+l_{3}\right)=m_{2}-k-m\geq m_{2}-\frac{m_{1}}{3}-\left(\frac{m_{1}}{6}+\frac{m_{2}}{2}\right)\geq 0\enspace.$ (3)
@@
    $Q^{*}\left(\mathsf{ESetInc}^{m_{1}+3m_{2}}_{m_{1}+m_{2},m_{1}+m_{2},m_{1},1/2}\right)$
    =$\quad$= $Q^{*}\left(\mathsf{ESetInc}^{4m+l}_{2m+l_{1}+l_{3},m+l_{2}+l_{3},k+l_{3},1/2}\right)$
[…]
    =$\quad$= $\Omega\left(\sqrt{m_{1}m_{2}}\right)\enspace.$
@@
$Q^{*}\left(\mathsf{ESetInc}^{n_{1}+3n_{2}}_{n_{1}+n_{2},n_{1}+n_{2},n_{1},g}\right)\geq Q^{*}\left(\mathsf{ESetInc}^{m_{1}+3m_{2}}_{m_{1}+m_{2},m_{1}+m_{2},m_{1},1/2}\right)=\Omega\left(\frac{\sqrt{n_{1}n_{2}}}{g}\right)\enspace.$
@@ 3.2 Randomized and Quantum Communication Complexity Upper Bound
Lemma 3.9 (Classical Upper Bound).
Fix $n,a,b\in\mathbb{Z}$. Consider $c,g$ such that $c+g,c-g\in\mathbb{Z}$. Let $n_{1}\coloneqq\min\{[a-c,c,b-c,n-a-b+c]\}$ and $n_{2}\coloneqq\min\left(\{[a-c,c,b-c,n-a-b+c]\}\setminus\left\{n_{1}\right\}\right)$. For any input $x,y\in\{0,1\}^{n}$ of $\mathsf{SetInc}_{a,b,c,g}^{n}$, there exists a randomized communication protocol that computes $\mathsf{SetInc}_{a,b,c,g}^{n}(x,y)$ using $O\left(\frac{n_{1}n_{2}}{g^{2}}\log n\log\log n\right)$ bits of communication with success probability at least $1-1/(6\log n)$.
@@
Fix $n,a,b\in\mathbb{Z}$. Consider $c,g$ such that $c+g,c-g\in\mathbb{Z}$. Let $n_{1}\coloneqq\min\{[a-c,c,b-c,n-a-b+c]\}$ and $n_{2}\coloneqq\min\left(\{[a-c,c,b-c,n-a-b+c]\}\setminus\left\{n_{1}\right\}\right)$. For any input $x,y\in\{0,1\}^{n}$ of $\mathsf{SetInc}_{a,b,c,g}^{n}$, there exists a quantum communication protocol without prior entanglement that computes $\mathsf{SetInc}_{a,b,c,g}^{n}(x,y)$ using $O\left(\frac{\sqrt{n_{1}n_{2}}}{g}\log n\log\log n\right)$ qubits of communication with success probability at least $1-1/(6\log n)$.
We note that Lemma 3.10 is a quantum speedup version of Lemma 3.9 by quantum amplitude amplification. The proof of Lemma 3.9 is given at the end of this section, and we postpone the proof of Lemma 3.10 to Section B.2.
@@
    Alice and Bob use binary search to determine $i\in\{0,1,...,k\}$ such that $\lvert x\land y\rvert\in I_{i}$, where $I_{i}$ is defined in Equation (4).
We first discuss the communication complexity of the above protocol. The first step takes $O(\log n)$ bits of communication. The fourth step costs $O\left(m(f)^{2}\log^{2}n\log\log n\right)$ bits of communication: For each $i\in[k]$, Alice and Bob can determine whether $|x\land y|\leq c_{i}-g_{i}$ or $|x\land y|\geq c_{i}+g_{i}$ by $O(m(f)^{2}\log n\log\log n)$ with a success probability of at least $1-1/\left(6\log n\right)$ by Lemma 3.9. Since binary search takes at most $\lceil\log\left(k+1\right)\rceil=O\left(\log n\right)$ rounds, the total communication cost is $O\left(m(f)^{2}\log^{2}n\log\log n\right)$.
@@
$\left\{I_{0}\coloneqq[0,c_{1}-g_{1}],I_{1}\coloneqq[c_{1}+g_{1},c_{2}-g_{2}],\dots,I_{k-1}\coloneqq[c_{k-1}+g_{k-1},c_{k}-g_{k}],I_{k}\coloneqq[c_{k}+g_{k},n]\right\}\enspace.$ (4)
@@
Therefore, Alice and Bob start from $i=\lfloor(k+1)/2\rfloor$ to determine whether $|x\land y|\leq c_{i}-g_{i}$ or $|x\land y|\geq c_{i}+g_{i}$ with success probability of at least $1-1/\left(6\log n\right)$. Depending on the result, they repeat the same process similar to binary search to find the interval that $|x\land y|$ falls in. After at most $\lceil\log\left(k+1\right)\rceil=O\left(\log n\right)$ repetitions, there is only one remaining interval and they can easily determine $f_{a,b}(|x\land y|)$. For $n\geq 2$, the failure probability of the above protocol is at most
$1-\left(1-\frac{1}{6\log n}\right)^{\lceil\log\left(k+1\right)\rceil}\leq\frac{\lceil\log\left(k+1\right)\rceil}{6\log n}\leq\frac{\lceil\log\left(n+1\right)\rceil}{3\log n^{2}}\leq\frac{1}{3}.$
@@
Fact 3.11 ([2, Lemma 30]).
Fix $10<\epsilon<\beta<$ 1 such that $\beta+\epsilon\leq 1$. For a set $S$, suppose there is a subset $S^{\prime}$ of $S$ such that $\frac{|S^{\prime}|}{|S|}\leq\beta-\epsilon$ or $\frac{|S^{\prime}|}{|S|}\geq\beta+\epsilon$. Suppose we can sample from $S$ uniformly and ask whether the sample is in $S^{\prime}$. Then we can decide whether $\frac{|S^{\prime}|}{|S|}\leq\beta-\epsilon$ or $\frac{|S^{\prime}|}{|S|}\geq\beta+\epsilon$ by $O(\beta/\epsilon^{2})$ samples, with success probability at least $2/3$.
Fact 3.12.
@@
We refer interesting readers to Section A.4 for the proof of 3.12. Now we prove the lemma by casing on the values of $n_{1}$ and $n_{2}$.
  • Case 1: $n_{1}=c$ and $n_{2}=a-c$. According to Definition 2.3, we have either $\frac{|x\land y|}{|x|}\leq\frac{c-g}{a}$ or $\frac{|x\land y|}{|x|}\geq\frac{c+g}{a}$. Alice and Bob estimate $\frac{|x\land y|}{|x|}$ as follows: Alice chooses an index $i$ such that $x_{i}=1$ uniformly at random. Then Alice sends $i$ to Bob, and Bob checks whether $y_{i}=1$. By 3.11, setting $\beta\coloneqq\frac{c}{a}$, $\epsilon\coloneqq\frac{g}{a}$, Alice and Bob can decide whether $\frac{|x\land y|}{|x|}\leq\frac{c-g}{a}$ or $\frac{|x\land y|}{|x|}\geq\frac{c+g}{a}$ with bounded error using $O\left(\frac{ac}{g^{2}}\right)=O\left(\frac{n_{1}n_{2}}{g^{2}}\right)$ samples. Since $|x|=a$, using $O\left(\frac{n_{1}n_{2}}{g^{2}}\log\log n\right)$ samples, they can decide whether $|x\land y|\leq c-g$ or $|x\land y|\geq c+g$ with success probability at least $1-1/(6\log n)$ by error reduction. Thus, the communication complexity is $O\left(\frac{n_{1}n_{2}}{g^{2}}\log n\log\log n\right)$.
@@
[…]

### v2

@@ 3.1 Quantum Communication Complexity Lower Bound
In this section, our goal is to obtain a lower bound on the quantum communication complexity for permutation-invariant functions (Theorem 3.1). Towards this end, we show that every permutation-invariant function $f$ can be reduced to $\mathrm{ESetInc}$ (defined in Definition 1.13) and exhibit a lower bound for $\mathrm{ESetInc}$ (Lemma 3.3). Additionally, Lemma 3.3 implies if $|x|=a,|y|=b$, then the cost to distinguish $|x\land y|=c-g$ from $|x\land y|=c+g$ is related to the smallest two numbers in $\{[a-c,c,b-c,n-a-b+c]\}$.
@@
Fix $n\in{\mathbb{Z^{+}}}$. Consider $a,b\in\left\{1,...,n-1\right\}$ and cg,c+gc-g,c+g are achievable Hamming weights of $|x\land y|$ when $|x|=a,|y|=b$. Let $n_{1}\coloneqq\min\{[a-c,c,b-c,n-a-b+c]\}$ and $n_{2}\coloneqq\min\left(\{[a-c,c,b-c,n-a-b+c]\}\setminus\left\{n_{1}\right\}\right)$. We have
$Q\left(\mathrm{ESetInc}_{a,b,c,g}^{n}\right)=\Omega\left(\frac{\sqrt{n_{1}n_{2}}}{g}\right).$
@@
By the definitions of $f_{a,b}$ and jump of $f_{a,b}$, any quantum protocol computing $f$ can also compute $\mathrm{ESetInc}^{n}_{a,b,c,g}$ for any $a,b$ and any jump $(c,g)\in\mathscr{J}(f_{a,b})$. Therefore, given a jump $(c,g)$ for $f_{a,b}$, the cost of computing $\mathrm{ESetInc}^{n}_{a,b,c,g}$ lower bounds the cost of computing $f$. By Lemma 3.3, we have $Q(f)\geq\Omega\left(\frac{\sqrt{n_{1}n_{2}}}{g}\right)$ for any jump $(c,g)$ in $f_{a,b}$, where $n_{1},n_{2}$ are the smallest two numbers in $\{[a-c,c,b-c,n-a-b+c]\}$. We conclude that $Q(f)=\Omega\left(m(f)\right)$ as desired. ∎
@@
Fix $n\in{\mathbb{Z^{+}}}$. Consider $a,b\in\left\{1,...,n-1\right\}$ and cg,c+gc-g,c+g are achievable Hamming weights of $|x\land y|$ when $|x|=a,|y|=b$. Let $n_{1}\coloneqq\min\{[a-c,c,b-c,n-a-b+c]\}$ and $n_{2}\coloneqq\min\left(\{[a-c,c,b-c,n-a-b+c]\}\setminus\left\{n_{1}\right\}\right)$. We have
$Q\left(\mathrm{ESetInc}_{a,b,c,g}^{n}\right)\geq Q\left(\mathrm{ESetInc}_{n_{1}+n_{2},n_{1}+n_{2},n_{1},g}^{n_{1}+3n_{2}}\right).$
@@
Consider $n_{1},n_{2},g$ such that $n_{1}\leq n_{2}$ and $n_{1}-g,n_{1}+g$ are achievable Hamming weights of $|x\land y|$ when $|x|=|y|=n_{1}+n_{2}$, we have
$Q\left(\mathrm{ESetInc}_{n_{1}+n_{2},n_{1}+n_{2},n_{1},g}^{n_{1}+3n_{2}}\right)=\Omega\left(\frac{\sqrt{n_{1}n_{2}}}{g}\right).$
We use the following two results on $\mathsf{ESetInc}$ to show Lemmas 3.4 and 3.5. Specifically, Lemma 3.6 is a variant of Lemma 4.1 in [17] and shows some reduction methods to the instances of the Exact Set-Inclusion Problem. Lemma 3.7 is a generalization of Theorem 5 in [8] proved by pattern matrix method and shows the lower bound of a special instance of the Exact Set-Inclusion Problem. The proofs of Lemmas 3.6 and 3.7 are given in Appendix B.1.
@@
Fix $n\in{\mathbb{Z^{+}}}$. Consider $a,b\in\left\{1,...,n-1\right\}$ and cg,c+gc-g,c+g are achievable Hamming weights of $|x\land y|$ when $|x|=a,|y|=b$. The following relations hold.
@@
    $Q\left(\mathrm{ESetInc}_{a,b,c,g}^{n}\right)\leq Q\left(\mathrm{ESetInc}_{a+\ell_{1}+\ell_{3},b+\ell_{2}+\ell_{3},c+\ell_{3},g}^{n+\ell}\right)$ for integers $\ell_{1},\ell_{2},\ell_{3}\geq 0$ such that $\ell_{1}+\ell_{2}+\ell_{3}\leq\ell$;
@@
    $Q\left(\mathrm{ESetInc}_{a,b,c,g}^{n}\right)=Q\left(\mathrm{ESetInc}_{a,n-b,a-c,g}^{n}\right)=Q\left(\mathrm{ESetInc}_{n-a,b,b-c,g}^{n}\right)$;
@@
    $Q\left(\mathrm{ESetInc}_{a,b,c,g}^{n}\right)\leq Q\left(\mathrm{ESetInc}_{ka,kb,kc,kg}^{kn}\right)$, where $k\geq 1$ is an integer.
@@
For every $k\in{\mathbb{Z}^{+}}$, if $l$ is a half-integer and $0<l\leq k/2$, then $Q\left(\mathrm{ESetInc}^{4k}_{2k,k,l,1/2}\right)=\Omega\left(\sqrt{kl}\right)$.
@@
$Q\left(\mathrm{ESetInc}^{n_{1}+3n_{2}}_{n_{1}+n_{2},n_{1}+n_{2},n_{1},g}\right)$
$=Q\left(\mathrm{ESetInc}^{n_{1}+n_{2}+n_{2}+n_{2}}_{n_{1}+n_{2},n_{1}+n_{2},n_{1},g}\right)$
$\leq Q\left(\mathrm{ESetInc}^{n_{1}+n_{2}+n_{3}+n_{4}}_{n_{1}+n_{2},n_{1}+n_{3},n_{1},g}\right)$
$=Q\left(\mathrm{ESetInc}^{n}_{a,b,c,g}\right).$
If $n_{2}=b-c$ or nab+cn-a-b+c, the argument is similar. ∎
@@
$Q\left(\mathrm{ESetInc}^{n_{1}+3n_{2}}_{n_{1}+n_{2},n_{1}+n_{2},n_{1},g}\right)$
$\geq Q\left(\mathrm{ESetInc}^{m_{1}+3m_{2}}_{m_{1}+m_{2},m_{1}+m_{2},m_{1},1/2}\right).$
@@
    $Q\left(\mathrm{ESetInc}^{m_{1}+3m_{2}}_{m_{1}+m_{2},m_{1}+m_{2},m_{1},1/2}\right)$
    $=\Omega\left(1\right)$
    $=\Omega\left(\sqrt{m_{1}m_{2}}\right).$
@@
    By LABEL:{le:SecInc} and 3.7, we have
    $Q\left(\mathrm{ESetInc}^{m_{1}+3m_{2}}_{m_{1}+m_{2},m_{1}+m_{2},m_{1},1/2}\right)$
    $=Q\left(\mathrm{ESetInc}^{4m^{\prime}_{2}+l}_{2m^{\prime}_{2}+l_{1},m^{\prime}_{2}+l_{2},m_{1},1/2}\right)$
    $\geq Q\left(\mathrm{ESetInc}^{4m^{\prime}_{2}}_{2m^{\prime}_{2},m^{\prime}_{2},m_{1},1/2}\right)$
@@
    $k\leq\frac{1}{2}\cdot\left\lfloor\frac{2m_{1}}{3}\right\rfloor\leq\frac{1}{2}\cdot\left\lfloor\frac{m_{1}}{6}+\frac{m_{2}}{2}\right\rfloor\leq\frac{m}{2},$ (2)
@@
    $l-\left(l_{1}+l_{2}+l_{3}\right)$ $=m_{2}-k-m$ (3)
    $\geq m_{2}-\frac{m_{1}}{3}-\left(\frac{m_{1}}{6}+\frac{m_{2}}{2}\right)$
    $\geq 0.$
@@
    $Q\left(\mathrm{ESetInc}^{m_{1}+3m_{2}}_{m_{1}+m_{2},m_{1}+m_{2},m_{1},1/2}\right)$
    $=Q\left(\mathrm{ESetInc}^{4m+l}_{2m+l_{1}+l_{3},m+l_{2}+l_{3},k+l_{3},1/2}\right)$
    $\geq Q\left(\mathrm{ESetInc}^{4m}_{2m,m,k,1/2}\right)$
    $=\Omega\left(\sqrt{mk}\right)$
    $=\Omega\left(\sqrt{m_{1}m_{2}}\right),$
    where the inequality follows from Lemma 3.6 and (3) and the second equality follows from Lemma 3.7 and (2).
@@
$Q\left(\mathrm{ESetInc}^{n_{1}+3n_{2}}_{n_{1}+n_{2},n_{1}+n_{2},n_{1},g}\right)$
$\geq Q\left(\mathrm{ESetInc}^{m_{1}+3m_{2}}_{m_{1}+m_{2},m_{1}+m_{2},m_{1},1/2}\right)$
$=\Omega\left(\frac{\sqrt{n_{1}n_{2}}}{g}\right).$
@@ 3.2 Randomized and Quantum Communication Complexity Upper Bound
The intuition of our randomized protocol to compute $\mathsf{SetInc}$ is as follows: Let $s_{1},s_{2}$ $(s_{1}\leq s_{2})$ be the smallest two numbers in $\{[|\overline{x}\land y|,|x\land y|,|\overline{x}\land y|,|\overline{x}\land\overline{y}|]\}$, and $n_{1},n_{2}$ $(n_{1}\leq n_{2})$ be the smallest two numbers in $\{[a-c,c,b-c,n-a-b+c]\}$. It is worth noting that $|n_{1}-s_{1}|=|n_{2}-s_{2}|=g$. Let $p=\frac{s_{1}}{s_{1}+s_{2}}$. Then we show $\mathsf{SetInc}$ is equivalent to distinguish $p\leq\beta-\epsilon$ from $p\geq\beta+\epsilon$ for some $\beta=O\left(\frac{n_{1}}{n_{2}}\right),\epsilon=\Omega\left(\frac{n_{2}}{g}\right)$. Depending on the value of $n_{1},n_{2}$, we use different sampling methods to estimate $p$ with error $\epsilon$. In the bounded-error randomized case, the communication cost is $O\left(\frac{\beta}{\epsilon^{2}}\log n\right)=O\left(\frac{n_{1}n_{2}}{g^{2}}\log n\right)$. Then we repeat $O(\log\log n)$ times of protocol such that the failed probability is at most $O(1/\log n)$. Finally, we use quantum amplitude amplification to speed up the randomized protocol.
@@
Lemma 3.8 (Randomized Upper Bound).
Fix $n,a,b\in{\mathbb{Z}^{+}}$. Consider $c,g$ such that $c+g,c-g\in{\mathbb{N}}$. Let $n_{1}\coloneqq\min\{[a-c,c,b-c,n-a-b+c]\}$ and $n_{2}\coloneqq\min\left(\{[a-c,c,b-c,n-a-b+c]\}\setminus\left\{n_{1}\right\}\right)$. For any input $x,y\in\{0,1\}^{n}$ of $\mathrm{SetInc}_{a,b,c,g}^{n}$, there exists a randomized communication protocol that computes $\mathrm{SetInc}_{a,b,c,g}^{n}(x,y)$ using $O\left(\frac{n_{1}n_{2}}{g^{2}}\log n\log\log n\right)$ bits of communication with success probability at least $1-1/(6\log n)$.
@@
Fix $n,a,b\in{\mathbb{Z}^{+}}$. Consider $c,g$ such that $c+g,c-g\in{\mathbb{N}}$. Let $n_{1}\coloneqq\min\{[a-c,c,b-c,n-a-b+c]\}$ and $n_{2}\coloneqq\min\left(\{[a-c,c,b-c,n-a-b+c]\}\setminus\left\{n_{1}\right\}\right)$. For any input $x,y\in\{0,1\}^{n}$ of $\mathrm{SetInc}_{a,b,c,g}^{n}$, there exists a quantum communication protocol without prior entanglement that computes $\mathrm{SetInc}_{a,b,c,g}^{n}(x,y)$ using $O\left(\frac{\sqrt{n_{1}n_{2}}}{g}\log n\log\log n\right)$ qubits of communication with success probability at least $1-1/(6\log n)$.
The proof of Lemma 3.8 is given at the end of this section, and we postpone the proof of Lemma 3.9 to Section B.2.
@@
    Alice and Bob use binary search to determine $i\in\{0,1,...,k\}$ such that $\lvert x\land y\rvert\in I_{i}$, where $I_{i}$ is defined in (4).
We first discuss the communication complexity of the above protocol. The first step takes $O(\log n)$ bits of communication. The fourth step costs $O\left(m(f)^{2}\log^{2}n\log\log n\right)$ bits of communication: For each $i\in[k]$, Alice and Bob can determine whether $|x\land y|\leq c_{i}-g_{i}$ or $|x\land y|\geq c_{i}+g_{i}$ by $O(m(f)^{2}\log n\log\log n)$ communication cost with a success probability of at least $1-1/\left(6\log n\right)$ by Lemma 3.8. Since binary search takes at most $\lceil\log\left(k+1\right)\rceil=O\left(\log n\right)$ rounds, the total communication cost is $O\left(m(f)^{2}\log^{2}n\log\log n+\log n\right)$.
@@
$I_{0}$ $\coloneqq[0,c_{1}-g_{1}],$ (4)
$I_{1}$ $\coloneqq[c_{1}+g_{1},c_{2}-g_{2}],$
$\dots$
$I_{k-1}$ $\coloneqq[c_{k-1}+g_{k-1},c_{k}-g_{k}],$
$I_{k}$ $\coloneqq[c_{k}+g_{k},n].$
@@
Therefore, Alice and Bob start from $i=\lfloor(k+1)/2\rfloor$ to determine whether $|x\land y|\leq c_{i}-g_{i}$ or $|x\land y|\geq c_{i}+g_{i}$ with success probability of at least $1-1/\left(6\log n\right)$. Depending on the result, they repeat the same process similar to binary search to find the interval that $|x\land y|$ falls in. After at most $\lceil\log\left(k+1\right)\rceil=O\left(\log n\right)$ repetitions, there is only one remaining interval and they can determine $f_{a,b}(|x\land y|)$. For $n\geq 2$, the failure probability of the above protocol is at most
$1-\left(1-\frac{1}{6\log n}\right)^{\lceil\log\left(k+1\right)\rceil}$ $\leq\frac{\lceil\log\left(k+1\right)\rceil}{6\log n}$
$\leq\frac{\lceil\log\left(n+1\right)\rceil}{3\log n^{2}}$
$\leq\frac{1}{3}.$
@@
Fact 3.10 ([1, Lemma 30]).
Fix $0<\epsilon<\beta<1$ such that $\beta+\epsilon\leq 1$. For a set $S$, suppose there is a subset $S^{\prime}$ of $S$ such that $\frac{|S^{\prime}|}{|S|}\leq\beta-\epsilon$ or $\frac{|S^{\prime}|}{|S|}\geq\beta+\epsilon$. Suppose we can sample from $S$ uniformly and ask whether the sample is in $S^{\prime}$. Then we can decide whether $\frac{|S^{\prime}|}{|S|}\leq\beta-\epsilon$ or $\frac{|S^{\prime}|}{|S|}\geq\beta+\epsilon$ by $O(\beta/\epsilon^{2})$ samples, with success probability at least $2/3$.
Fact 3.11.
@@
We refer interesting readers to Section A.4 for the proof of 3.11. Now we prove the lemma by casing on the values of $n_{1}$ and $n_{2}$.
  • Case 1: $n_{1}=c$ and $n_{2}=a-c$. According to Definition 1.13, we have either $\frac{|x\land y|}{|x|}\leq\frac{c-g}{a}$ or $\frac{|x\land y|}{|x|}\geq\frac{c+g}{a}$. Alice and Bob estimate $\frac{|x\land y|}{|x|}$ as follows: Alice chooses an index $i$ such that $x_{i}=1$ uniformly at random. Then Alice sends $i$ to Bob, and Bob checks whether $y_{i}=1$. By 3.10, setting $\beta\coloneqq\frac{c}{a}$, $\epsilon\coloneqq\frac{g}{a}$, Alice and Bob can decide whether $\frac{|x\land y|}{|x|}\leq\frac{c-g}{a}$ or $\frac{|x\land y|}{|x|}\geq\frac{c+g}{a}$ with bounded error using $O\left(\frac{ac}{g^{2}}\right)=O\left(\frac{n_{1}n_{2}}{g^{2}}\right)$ samples. Since $|x|=a$, using $O\left(\frac{n_{1}n_{2}}{g^{2}}\log\log n\right)$ samples, they can decide whether $|x\land y|\leq c-g$ or $|x\land y|\geq c+g$ with success probability at least $1-1/(6\log n)$ by error reduction. Thus, the communication complexity is $O\left(\frac{n_{1}n_{2}}{g^{2}}\log n\log\log n\right)$.
@@
      $|\overline{x}\oplus y|$ $=|x\land y|+|\overline{x}\land\overline{y}|$
      $=|x\land y|+\left(n-\left(a+b-|x\land y|\right)\right)$
      $=n-(a+b)+2|x\land y|,$